OpenStudy (anonymous):

Simplify without negative exponents. x^-4 / x^7

OpenStudy (anonymous):

\[\frac{x^{-4}}{x^7}=\frac{1}{x^4x^7}=\frac{1}{x^{4+7}}\] etc